In a right triangle, sin θ = 9/15. Find the value of tan θ
Answer by ewatrrr(24785)

 
Hi,
In a right triangle, sin θ = 9/15 = opp/hyp
15^2 - 9^2 = 144, 12 is the adj side
tanθ = opp/adj = 9/12 = 3/4